# Computing the action of a matrix exponential on an interval via the $\star$-product approach

### v1

@@ Abstract
We present a new method for computing the action of the matrix exponential on a vector, $e^{At}v$. The proposed approach efficiently evaluates the solution for all $t$ within a prescribed bounded interval by expanding it into an orthogonal polynomial series. This method is derived from a new representation of the matrix exponential in the so-called $\star$-algebra, an algebra of bivariate distributions. The resulting formulation leads to a linear system equivalent to a matrix equation of Stein type, which can be solved by either direct or Krylov subspace methods. Numerical experiments demonstrate the accuracy and efficiency of the proposed approach in comparison with state-of-the-art techniques.
@@ 1 Introduction
The approach proposed in this work builds on the recently developed theory of $\star$-algebra [16, 15, 17]. By discretizing the $\star$-product via Legendre polynomial expansions, it reformulates the problem of matrix approximation into a linear algebra problem, which can be equivalently expressed in terms of a Stein-type matrix equation. The $\star$-approach gives the solution in terms of the coefficient of the truncated Legendre polynomial expansion of $e^{tA}v$ in a given interval $\mathcal{I}$. Therefore, the solution can be computed for every $t\in\mathcal{I}$. This paper also provides a bound for the truncation error of the method. Combined with an Arnoldi-based Krylov projection of dimension $k$, the method yields an efficient algorithm for large sparse matrices, allowing the solution to be evaluated at arbitrary times without recomputation.
@@ 2 Mathematical Formulation
with $\delta_{k\ell}$ the Kronecker delta. Since these polynomials form a basis for analytic functions over $\mathcal{I}$, for a given positive integer $M$, we can approximate each entry of the solution $U(t,s)$ by a truncated double expansion.
@@
where $I_{N}$ is the identity matrix of size $N$, and $\mathcal{U}_{M}$ is named the coefficient matrix of $U(t,s)$ and defined as the block matrix
@@
Substituting approximation (7) into approximation (5) yields
@@
with exact solution $\tilde{u}(t)=\exp\!\big(\frac{1}{2}A(t+1)\big)v$. Note that, $\tilde{u}(t)=\exp\!\big(\frac{1}{2}A(t+1)\big)v=\exp\!\big(A\,\hat{t}\big)v$, for $t\in[-1,1],\,\hat{t}\in[0,1]$. Using Eq. (8), the solution of Eq.(9) is approximated by
@@ 4.1 Algorithms
The first method to solve Eq. (9) is based on the Schur decomposition of the coefficient matrix associated with $\Theta(t-s)$, i.e., $T_{M}=USU^{H},$ where $U$ is unitary and $S$ is a upper triangular matrix. With the change of variables $Y=U^{H}X$ and $w=U^{H}\phi_{M}(-1)$, Eq. (11) becomes
@@
The Schur decomposition is performed only once, stored, and reused; the overall computational cost is therefore dominated by solving (21). We consider two variants for computing the solution:

### v2

@@ Abstract
We present a new method for computing the action of the matrix exponential on a vector, $e^{At}v$. The proposed approach efficiently evaluates the solution for all $t$ within a prescribed bounded interval by expanding it into an orthogonal polynomial series. This method is derived from a new representation of the matrix exponential in the so-called $\star$-algebra, an algebra of bivariate distributions. The resulting formulation leads to a linear system equivalent to a matrix equation of Stein type, which can be solved by either direct or Krylov subspace methods. Numerical experiments demonstrate the accuracy and efficiency of the proposed approach in comparison to state-of-the-art techniques.
@@ 1 Introduction
The approach proposed in this work is based on the recently developed theory of $\star$-algebra [16, 15, 17]. By discretizing the $\star$-product via Legendre polynomial expansions, it reformulates the problem of matrix approximation into a linear algebra problem, which can be equivalently expressed in terms of a Stein-type matrix equation. The $\star$-approach gives the solution in terms of the coefficient of the truncated Legendre polynomial expansion of $e^{tA}v$ in a given interval $\mathcal{I}$. Therefore, the solution can be computed for every $t\in\mathcal{I}$. This paper also provides a bound for the truncation error of the method. Combined with an Arnoldi-based Krylov projection of dimension $k$, the method yields an efficient algorithm for large sparse matrices, allowing the solution to be evaluated at arbitrary times without recomputation.
@@ 2 Mathematical Formulation
with $\delta_{k\ell}$ the Kronecker delta. Since these polynomials form the basis for the analytic functions in $\mathcal{I}$, for a given positive integer $M$, we can approximate each entry of the solution $U(t,s)$ by a truncated double expansion.
@@
where $I_{N}$ is the identity matrix of size $N$, and $\mathcal{U}_{M}$ is named the coefficient matrix of $U(t,s)$ and is defined as the block matrix
@@
Substituting the approximation (7) into approximation (5) yields
@@
with an exact solution $\tilde{u}(t)=\exp\!\big(\frac{1}{2}A(t+1)\big)v$. Note that $\tilde{u}(t)=\exp\!\big(\frac{1}{2}A(t+1)\big)v=\exp\!\big(A\,\hat{t}\big)v$, for $t\in[-1,1],\,\hat{t}\in[0,1]$.
Using Eq. (8), the solution of Eq.(9) is approximated by
@@ 4.1 Algorithms
The first method to solve Eq. (9) is based on the Schur’s decomposition of the coefficient matrix associated with $\Theta(t-s)$, i.e., $T_{M}=USU^{H},$ where $U$ is unitary and $S$ is a upper triangular matrix. With the change of variables $Y=U^{H}X$ and $w=U^{H}\phi_{M}(-1)$, Eq. (11) becomes
@@
Schur decomposition is performed only once, stored, and reused; the overall computational cost is therefore dominated by solving (21). We consider two variants for computing the solution:
@@ Acknowledgments
Declaration of generative AI and AI-assisted technologies
OpenAI’s ChatGPT (GPT‑5) and Microsoft 365 Copilot (bizchat.20260210.47.1) were used for language polishing and for assisting with bibliographic research, as well as with trivial mathematical derivations, all of which were strictly supervised and verified by the authors.